Blevins Boards produces two kinds of skateboards. Selected unit data for the two boards for the last quarter follow:
 
Blevins allocates production overhead using activity-based costing. It allocates delivery expense and sales commissions, which amount to $54,000 per quarter, to the two products equally.
 
Required
a. Compute the per-unit cost for each product.
b. Compute the profit for each product.
Answer #1
Delivery expense & Sales comission $54000 per quarter
Equal distribution indicates $27000 to each product
Basco boards per unit cost = $27000/5000 = $5.4 per unit
Shimano boards per unit cost = $27000/10000 = $2.7
BASCO BOARDS SHIMANO BOARDS
Diret material 25 38
Direct labor 32 54
Allocated overheads 15 18
Delivery expense & Sales comission 5.4 2.7
a) Per unit cost 77.4 112.7
b) Units sold 5000 10000
c) Total cost 387000 1127000
d) Sales revenue 500000 1400000
e)= d-c Profit for each product per quarter 113000 273000